A sample of O2 gas occupies a volume of 1.50 L at a pressure of 735 mm Hg and a temperature of 25oC. What volume will it occupy, in liters, if the pressure is increased to 770 mm Hg with no change in temperature?
1
Expert's answer
2021-09-20T02:25:08-0400

P1V1/T1 = P2V2/T2

735 Γ— 1.5/ 298 = 770Γ— V2/298

3.7 = 2.58V2

V2 = 3.7/2.58

V2 = 1.43L